### v2

# concert_supp

Abstract
The popularity of transfer learning stems from the fact that it can borrow information from useful auxiliary datasets. Existing statistical transfer learning methods usually adopt a global similarity measure between the source data and the target data, which may lead to inefficiency when only partial information is shared. In this paper, we propose a novel Bayesian transfer learning method named “CONCERT” to allow robust partial information transfer for high-dimensional data analysis. A conditional spike-and-slab prior is introduced in the joint distribution of target and source parameters for information transfer. By incorporating covariate-specific priors, we can characterize partial similarities and integrate source information collaboratively to improve the performance on the target. In contrast to existing work, the CONCERT is a one-step procedure, which achieves variable selection and information transfer simultaneously. We establish variable selection consistency, as well as estimation and prediction error bounds for CONCERT. Our theory demonstrates the covariate-specific benefit of transfer learning. To ensure that our algorithm is scalable, we adopt the variational Bayes framework to facilitate implementation. Extensive experiments and two real data applications showcase the validity and advantage of CONCERT over existing cutting-edge transfer learning methods.
Keywords: Data heterogeneity, high-dimensional data, similarity selection, sparsity, variational Bayes
1 Introduction
The growing emphasis on data heterogeneity in the era of big data has brought high demand for developing efficient and robust methods to integrate information from heterogeneous data sources. The homogeneous data assumption, which is commonly assumed in the traditional machine learning and statistical literature, can lead to biased inference. As a remedy, transfer learning (Pan and Yang, 2009) has emerged for its ability to borrow information from similar but different tasks, which can enhance the estimation efficiency on the target data. This is especially useful for high-dimensional data analysis, where the number of subjects in the target dataset could be much less than the number of covariates. There is a rapidly growing literature on the application of transfer learning methods in domain science, including human activity recognition (Hirooka et al., 2022), protein representation learning (Fenoy et al., 2022), and RNA modification identification (Wu et al., 2024).
One of the challenges in transfer learning lies in how to extract useful information from source datasets, and statistical research on transfer learning has grown rapidly in recent years. Among them, Li et al. (2022), Tian and Feng (2023) and Li et al. (2023) focused on (generalized) linear models, Li et al. (2023) investigated Gaussian graphical models, Zhang and Zhu (2022) studied quantile models, and Lin and Reimherr (2022) explored functional data. The notion of these methods is to measure the similarity between the source and target based on the global difference between parameters and adaptively choose informative sources with small global differences. However, the global similarity measure may lead to an all-in-or-all-out principle where each source dataset is either fully utilized for information transfer or completely disregarded, resulting in inefficiency.
To see when global similarity may fail, we illustrate two cases of partial similarity structure in the left panel of Figure 1. Figure 1(a) represents the case where sources contain heterogeneous information and only share part of the parameters with the target, and Figure 1(c) demonstrates the case where sources contain all significant signals of the target but have redundant information on additional dimensions.
Figure 1: Illustration of the two cases of partial information similarity. In (a) and (c), the true variable identities are clarified with Target Positive, Source Positive, and Negative representing influential variables in the target, transferable variables in sources, and negative variables, respectively. In (b) and (d), the corresponding estimated posterior probabilities of all variables are shown in terms of significance in the target and transferability in sources.
Both cases are common in image classification or medical data. However, due to the global similarity requirement, existing methods may fail to recognize the sources as informative auxiliary data even if they share some partial structure. This motivates us to develop new similarity measures to fully utilize the information from sources in the absence of global similarity.
Another challenge arises from the unknown similarities between the target and source datasets. To avoid negative transfer (Zhang et al., 2023), a two-stage framework has been developed in the transfer learning literature (Li et al., 2022; Tian and Feng, 2023; Li et al., 2023), where the global similarity between sources and the target is evaluated first. Informative sources are then selected based on case-specific criteria, or a model averaging procedure is implemented. Since the underlying similarity is not known in advance, selection methods usually rely on a global similarity gap assumption between informative and non-informative sources (Tian and Feng, 2023; Zhang and Zhu, 2022) to ensure consistent source selection. In the absence of global similarity, detecting similarity becomes more challenging. Therefore, a partially adaptive and robust transfer procedure is needed.
Motivated by the above challenges, we propose a novel Bayesian transfer learning method designed to enable robust partial information transfer for high-dimensional generalized linear regression. We introduce a new conditional spike-and-slab prior in conjunction with the traditional spike-and-slab prior (George and McCulloch, 1993; Ishwaran and Rao, 2005; Ročková and George, 2018) for the joint distribution of parameters across both the target and sources. To characterize the partial similarity, we equip the coefficient of each covariate from each source with a conditional spike-and-slab prior centered at the corresponding target parameter. With this joint prior, our method can adaptively achieve variable selection and information transfer in a one-step manner. We name our proposed method CONCERT for CONditional spike-and-slab for Covariate-Elaborated Robust Transfer, which robustly transfers information from informative covariates across different sources, ensuring they can be collaboratively integrated to improve the target. In Figure 1(b) and 1(d), we show the validity of CONCERT on joint variable and similarity selection. Details can be found in Section LABEL:append:simu1 of Supplementary Material.
Several independent lines of work provide reasons for Bayesian transfer learning. Among them the target prior is assigned conditionally based on source knowledge, either directly on source data (Karbalayghareh et al., 2018) or on source predictive density (Papež and Quinn, 2022; Jirsa et al., 2021), and the conditional posterior of target response is under concern. Our method differs significantly from these approaches by introducing a joint association of the parameters rather than directly assuming a joint distribution on the target and source data. The proposed method can significantly reduce model complexity while enhancing both efficiency and interpretability.
The CONCERT contributes to the literature in four aspects. Firstly, we allow partial information transfer. By characterizing the covariate-specific similarity for each source using the conditional spike-and-slab prior, we can fully utilize the partially useful information from different sources even in the absence of global similarity. We showcase the advantage over existing transfer learning methods through extensive numerical experiments and real data analysis. Secondly, we establish the variable selection consistency. Specifically, we relax the minimal signal strength condition in the traditional high-dimensional variable selection literature in a covariate-specific manner. Our theory demonstrates that the more similar sources with the target on a particular variable, the lower the minimum signal strength required to detect that variable in the target. In other words, we can detect weaker signals in the target with the benefit of borrowing partial information from more sources. The estimation and prediction errors of the CONCERT also enjoy faster convergence rates than single-task Bayesian regression.
Thirdly, the adoption of spike-and-slab prior for transfer learning here also provides a new perspective beyond the scope of conventional spike-and-slab prior, which is initially designed for variable selection. Under a mild local similarity gap condition, we also establish the consistency of source similarity detection under the conditional spike-and-slab prior. Lastly, we utilize variational Bayes to make our algorithm scalable with an ignorable loss of accuracy. Beyond the common mean-field variational family, a novel hierarchical structure is adopted to incorporate parameter dependence.
The rest of the paper is organized as follows. In Section 2, we introduce the problem setup and propose our CONCERT method for partial information transfer. A scalable variational Bayes algorithm is presented in Section 3. Theoretical guarantees for our method are provided in Section 4. Extensive experiments and two real data applications are provided in Sections 5 and 6, respectively. We conclude this paper in Section 7.
2 Methodology
In this section, we will introduce our CONCERT method. Consider the high-dimensional generalized linear models with response $y\in\mathbb{R}$ and covariates $x\in\mathbb{R}^{p}$. The conditional distribution of $y$ given $x$ follows
$p(y\mid x)=\varphi(y)\exp\{yx^{\top}\boldsymbol{\beta}-\psi(x^{\top}\boldsymbol{\beta})\},$ (1)
where $\boldsymbol{\beta}\in\mathbb{R}^{p}$ is the parameter, and $\varphi$ and $\psi$ are known functions with $\psi^{\prime}(x^{\top}\boldsymbol{\beta})=\mathbb{E}[y\mid x]$. Note that model (1) with $\psi(u)=u^{2}/2$ corresponds to the Gaussian linear regression model, while $\psi(u)=\log(1+e^{u})$ corresponds to the logistic regression model.
Suppose we are interested in the target data ${\cal D}^{(0)}=\{(x^{(0)}_{i},y^{(0)}_{i})\}^{n_{0}}_{i=1}$, where the conditional distribution satisfies (1):
$p(y_{i}^{(0)}\mid x^{(0)}_{i})=\varphi(y_{i}^{(0)})\exp\{y_{i}^{(0)}x^{(0)\top}_{i}\boldsymbol{\beta}^{(0)}-\psi(x^{(0)\top}_{i}\boldsymbol{\beta}^{(0)})\}.$
However, the sample size of the target data could be insufficient under the high-dimensional setting where $n_{0}\ll p$. To this end, we can utilize additional auxiliary datasets. Suppose there are $K$ additional independent source datasets ${\cal D}^{(k)}=\{(x^{(k)}_{i},y^{(k)}_{i})\}^{n_{k}}_{i=1}$ satisfying
$p(y_{i}^{(k)}\mid x^{(k)}_{i})=\varphi(y^{(k)}_{i})\exp\{y_{i}^{(k)}x^{(k)\top}_{i}\boldsymbol{\beta}^{(k)}-\psi(x^{(k)\top}_{i}\boldsymbol{\beta}^{(k)})\},$
for $k=1,\ldots,K$, respectively. Although additional information is provided by the samples of these source datasets, the heterogeneity between the source and the target may hinder the performance of existing methods based solely on the target data. Our aim is to extract as much useful information as possible from the source data to help improve the performance on the target while preventing negative transfer.
2.1 Spike-and-Slab for Variable Selection
Before we introduce our tailored transfer learning framework, we first provide more background on spike-and-slab for variable selection. In the Bayesian framework, the spike-and-slab prior has been widely adopted to perform variable selection, where a mixture of distributions is provided as the prior for the interested parameters. A crucial advantage of spike-and-slab is its availability of quantifying uncertainty for both parameter estimation and model selection (Narisetty et al., 2019; Nie and Ročková, 2023). Various mixture strategies have been adopted, including the absolute continuous spike-and-slab (Biswas et al., 2022) and the point-mass spike (Zeng et al., 2022). In this paper, we choose the Gaussian distribution as the slab prior and the Dirac mass at zero as the spike prior. If we only had the target data, let the priors on $\boldsymbol{\beta}^{(0)}$ be
$\beta^{(0)}_{j}\mid Z_{j}\stackrel{{\scriptstyle ind}}{{\sim}}Z_{j}N(0,\eta^{2})+(1-Z_{j})\delta_{0},\quad Z_{j}\stackrel{{\scriptstyle iid}}{{\sim}}\operatorname{Bernoulli}(q_{0}),$
where $Z_{j}$’s are the latent variable selection indicators showing whether the $j\text{th}$ covariate of $x^{(0)}$ is informative, i.e., whether its corresponding coefficient $\beta^{(0)}_{j}$ is away from zero for $j=1,\ldots,p$, $\delta_{x}$ denotes the Dirac mass at $x$, and $\eta$ and $q_{0}$ are the hyperparameters. Based on the likelihood and priors, one can make inferences from the posterior distribution on both variable selection and parameter estimation.
2.2 Conditional Spike-and-Slab for Similarity Selection
Now we illustrate how spike-and-slab priors can also be used to transfer information from sources to the target in a covariate-specific manner. A source data is defined to be helpful for estimating the target model in the sense that the information carried by a subset of covariates is similar to that in the target data. One simple example is $\boldsymbol{\beta}^{(k)}\approx\boldsymbol{\beta}^{(0)}$ where the entire set of covariates is helpful and can be transferred. However, a more reasonable case is that only a fraction of the covariates are useful, that is, $\boldsymbol{\beta}^{(k)}_{S}\approx\boldsymbol{\beta}^{(0)}_{S}$ for some $S\subseteq\{1,\ldots,p\}$ and other covariates are not informative. Though such source data is not fully useful, we can still selectively exploit relevant information to improve target learning.
Motivated by this purpose, we design a novel conditional spike-and-slab prior to select the similarity structure:
$\beta^{(k)}_{j}\mid\beta^{(0)}_{j},I^{(k)}_{j}\stackrel{{\scriptstyle ind}}{{\sim}}I^{(k)}_{j}N(\beta^{(0)}_{j},\tau_{k}^{2})+(1-I^{(k)}_{j})\delta_{\beta^{(0)}_{j}},\quad I^{(k)}_{j}\stackrel{{\scriptstyle iid}}{{\sim}}\operatorname{Bernoulli}(q_{k}),$
where $I^{(k)}_{j}$’s are the latent covariate-specific similarity indicators showing whether the $j\text{th}$ covariate of source $k$ is transferable for $j=1,\ldots,p$ and $k=1,\ldots,K$, and $\tau_{k}$ and $q_{k}$ are the hyperparameters. Our idea differs from the conventional spike-and-slab prior designed for variable selection. To better interpret our conditional spike-and-slab prior, we can define $\delta^{(k)}_{j}=\beta^{(k)}_{j}-\beta^{(0)}_{j}$ and rewrite the prior as
$\delta^{(k)}_{j}\mid\beta^{(0)}_{j},I^{(k)}_{j}\stackrel{{\scriptstyle ind}}{{\sim}}I^{(k)}_{j}N(0,\tau_{k}^{2})+(1-I^{(k)}_{j})\delta_{0},$
which inherits the spirit of mixture modeling in spike-and-slab. Notably different from the traditional spike-and-slab, the conditional spike-and-slab is designed for similarity selection, and the indicators take the role of selecting informative covariates from different sources. If $I^{(k)}_{j}=0$, the prior of $\beta^{(k)}_{j}$ concentrates on the point mass at $\beta^{(0)}_{j}$, showing a prior belief on $\beta^{(k)}_{j}=\beta^{(0)}_{j}$ and thus the corresponding covariate being transferable. Otherwise, if $I^{(k)}_{j}=1$, the prior of $\beta^{(k)}_{j}$ is diffusive away from $\beta^{(0)}_{j}$ under a large variance $\tau^{2}_{k}$, and the corresponding covariate is not considered useful for the target data.
Through the covariate-specific indicator, we can explore the partial similarity between the source and the target. Existing transfer learning methods (Li et al., 2022; Lin and Reimherr, 2022; Tian and Feng, 2023; Li et al., 2023) usually define informative sources based on the norm of difference $\boldsymbol{\beta}^{(k)}-\boldsymbol{\beta}^{(0)}$. This global similarity measure may fail to fully utilize useful information from partially similar sources, as illustrated in Figure 1. Here we adopt a new partial similarity measure to fill this gap. As we will see in our numerical studies in Section 5, our method outperforms existing methods in both the globally informative setting and the partially informative setting.
Remark 1.
In contrast to previous works in Bayesian transfer learning (Karbalayghareh et al., 2018; Papež and Quinn, 2022), our method does not impose additional assumptions on the target prior. Our target prior does not involve source information and thus remains in the original form of the sole target model, which blesses it with model consistency and interpretability despite the existence of sources.
3 Variational Bayes Implementation
The proposed Bayesian transfer learning can be carried out using standard Gibbs sampling from full conditional posterior distributions. However, here we adopt a scalable approach based on variational Bayes (Ray et al., 2020; Ray and Szabó, 2022). Variational Bayes (VB) serves as a popular tool to improve Bayesian methods on the computational cost of MCMC (Blei et al., 2017). The thrust of VB is to approximate the joint posterior distribution by the closest distribution within some tractable distribution family. VB is commonly used to assist stochastic search variable selection for spike-and-slab prior (Ray et al., 2020; Menacher et al., 2023), where its performance is guaranteed both empirically and theoretically (Carbonetto and Stephens, 2012; Ray and Szabó, 2022). In this section, we derive the VB algorithm for transfer learning.
We denote the likelihood functions of the target data and the source data as $L({\cal D}^{(0)})$ and $L({\cal D}^{(k)})$ for $k=1,\ldots,K$, respectively. Let $\theta^{(0)}$ and $\theta^{(k)}$ for $k=1,\ldots,K$ denote all parameters in the target and source models and $\Pi(\theta^{(0)})$ and $\Pi(\theta^{(k)})$ be the corresponding prior distributions. The joint prior distribution is thus given by $\Pi(\theta)=\bigotimes^{K}_{k=0}\Pi(\theta^{(k)})$. Let $\cal{D}$ denote all data. By Bayes’ formula the joint posterior density can be expressed as
$d\Pi(\theta\mid{\cal D})\propto L({\cal D}^{(0)})\prod^{K}_{k=1}L({\cal D}^{(k)})d\Pi(\theta).$ (2)
We propose a hierarchical mean-field variational family $\cal Q$ where $\beta^{(0)}_{j}$’s follow independent spike-and-slab distributions for $j=1,\ldots,p$:
$\beta^{(0)}_{j}\mid Z_{j}\sim Z_{j}N(\mu^{(0)}_{j},(\sigma^{(0)}_{j})^{2})+(1-Z_{j})\delta_{0},\quad Z_{j}\sim\operatorname{Bernoulli}(\gamma^{(0)}_{j}),$
and $\beta^{(k)}_{j}$’s are independent from each other while dependent on $\beta^{(0)}_{j}$ for $k=1,\ldots,K$:
$\beta^{(k)}_{j}\mid\beta^{(0)}_{j},I^{(k)}_{j}\sim I^{(k)}_{j}N(\mu^{(k)}_{j},(\sigma^{(k)}_{j})^{2})+(1-I^{(k)}_{j})\delta_{\beta^{(0)}_{j}},\quad I^{(k)}_{j}\sim\operatorname{Bernoulli}(\gamma^{(k)}_{j}).$
We further denote the variational distribution of other parameters belonging to each specific model as $Q(\phi^{(k)})$ for $k=0,\ldots,K$. Then $\cal Q$ can be represented as
$\mathcal{Q}=\left\{Q(\theta)=\bigotimes_{k=0}^{K}\left[\bigotimes_{j=1}^{p}\left(\gamma^{(k)}_{j}N(\mu^{(k)}_{j},(\sigma^{(k)}_{j})^{2})+(1-\gamma^{(k)}_{j})\Delta^{(k)}_{j}\right)\times Q(\phi^{(k)})\right]\right\},$ (3)
where $\Delta^{(k)}_{j}$ denotes the corresponding Dirac mass distributions. The variational posterior $Q^{*}$ is the element of $\cal Q$ which minimizes the KL divergence to the exact posterior distribution, or equivalently maximizes the evidence lower bound (ELBO):
$Q^{*}=\operatorname{argmax}_{Q\in{\cal Q}}\operatorname{ELBO}\equiv\operatorname{argmax}_{Q\in{\cal Q}}\mathbb{E}_{Q}\left[\log\frac{d\Pi(\theta\mid{\cal D})}{dQ(\theta)}\right].$ (4)
While the exact posterior distribution contains a sum of $2^{(K+1)p}$ models, the VB family $\cal Q$ replaces them with only $(K+1)p$ marginal models, drastically reducing the computational burden in the high-dimensional case.
3.1 Example: Linear Regression
Under the linear regression setting, the target and source models satisfy $y_{i}^{(k)}=x^{(k)\top}_{i}\boldsymbol{\beta}^{(k)}+\varepsilon_{i}^{(k)}$, where $\varepsilon_{i}^{(k)}$’s are i.i.d. random noises from $N(0,(\sigma^{(k)}_{y})^{2})$ for $k=0,\ldots,K$, respectively. We choose a commonly used inverse-Gamma prior distribution $\operatorname{I\Gamma}(a_{0},b_{0})$ on the additional parameters $\{(\sigma^{(k)}_{y})^{2}\}^{K}_{k=0}$, and assume the prior slab variances of the target and sources contain $(\sigma^{(k)}_{y})^{2}$ in addition to $\eta$ and $\tau_{k}$ for conjugacy, respectively. The hyperparameters in the weakly informative inverse-Gamma prior distributions are set as $a_{0}=2$ and $b_{0}=1$ in our numerical studies. For the VB implementation, the variational distributions for $(\sigma^{(k)}_{y})^{2}$ in (3) are assumed to be inverse-Gamma distributions $\operatorname{I\Gamma}(a^{(k)},b^{(k)})$ for $k=0,\ldots,K$.
Since the assigned prior distributions are conditionally conjugate in the linear regression scenario, coordinate ascent variational inference (CAVI) (Bishop, 2006) can be used to solve (4), which iteratively optimizes each factorized term while keeping the others fixed. We defer explicit updates and implementation details to Supplementary Material Section LABEL:supp:algo.
3.2 Example: Logistic Regression
Under the logistic regression setting, the target and source models satisfy
$P(y_{i}^{(k)}=1\mid x_{i}^{(k)})=\frac{\exp\{x^{(k)\top}_{i}\boldsymbol{\beta}^{(k)}\}}{1+\exp\{x^{(k)\top}_{i}\boldsymbol{\beta}^{(k)}\}}.$
Unlike linear regression models, the logistic regression models do not enjoy direct conjugacy between likelihood and spike-and-slab priors, which makes MCMC and VB schemes intractable. One way to overcome this difficulty is to adopt Pólya-Gamma (PG) data augmentation (Polson et al., 2013). We introduce a latent PG variable $\omega^{(k)}_{i}$ for each binary response $y^{(k)}_{i}$ and therefore the augmented likelihood $L({\cal D}\mid\omega)$ is given by
$L({\cal D}\mid\omega)=\prod^{K}_{k=0}\prod^{n_{k}}_{i=1}\frac{1}{2}\exp\left\{\left(y^{(k)}_{i}-\frac{1}{2}\right)x^{(k)\top}_{i}\boldsymbol{\beta}^{(k)}-\frac{1}{2}\omega^{(k)}_{i}(x^{(k)\top}_{i}\boldsymbol{\beta}^{(k)})^{2}\right\},$
where $\omega=(\omega^{(0)}_{1},\ldots,\omega^{(K)}_{n_{K}})$ with each $\omega^{(k)}_{i}$ following the PG distribution $\operatorname{PG}(1,0)$ for $i=1,\ldots,n_{k}$ and $k=0,\ldots,K$. The corresponding variational distributions of $\omega^{(k)}_{i}$’s are assumed to be $\operatorname{PG}(1,c^{(k)}_{i})$. Data augmentation guarantees conjugacy so that CAVI can be utilized to solve (4). A more detailed derivation can be found in Supplementary Material Section LABEL:supp:algo.
4 Theoretical Results
In this section, we study the theoretical properties of CONCERT under the linear regression setting. Without loss of generality, we assume that $\sigma_{y}^{(0)}=\cdots=\sigma_{y}^{(K)}$ is known in our theoretical analysis. All the proofs can be found in Supplementary Material Section LABEL:app:theorem.
We first introduce some notations. Let $[K]=\{1,\ldots,K\}$ and $N=\sum_{k=0}^{K}n_{k}$. Let $S=\{j\in[p]:\beta_{j}^{(0)}\neq 0\}$ be the target signal set. We denote $T_{k}=\{j\in[p]:\delta_{j}^{(k)}\neq 0\}$ to be the non-transferable index set for source $k$, and $T=\{T_{k}\}_{k=1}^{K}$. We consider the models with $|S|\leq L_{s}$ and $|T_{k}|\leq L_{t}$ for some large numbers $L_{s}$ and $L_{t}$. Let $\{\boldsymbol{\beta}^{(0)*}$, $\{\boldsymbol{\delta}^{(k)*}\}_{k=1}^{K}\}$ be the true parameters. We define the true signal and non-transferable index sets as
$S^{*}=\{j\in[p]:|\beta_{j}^{(0)*}|\geq\zeta_{j}\}\quad\text{and}\quad T^{*}_{k}=\{j\in[p]:\delta_{j}^{(k)*}\neq 0\},$ (5)
respectively, with $s^{*}=|S^{*}|$ and $t_{k}^{*}=|T_{k}^{*}|$ denoting the true target model size and true non-transferable model size of source $k$ for $k\in[K]$.
For a matrix $A$ and an index set $D$, let the sub-matrix of $A$ with columns in $D$ as $A_{D}$. Denote the design matrix on the $k$th dataset as $X^{(k)}$. Then $\tilde{P}_{T_{k}}=X^{(k)}_{T_{k}}(X^{(k)\top}_{T_{k}}X^{(k)}_{T_{k}})^{-1}X^{(k)\top}_{T_{k}}$ and $\tilde{Q}_{T_{k}}=I_{n_{k}}-\tilde{P}_{T_{k}}$ are the projection matrices on the column space of $X_{T_{k}}^{(k)}$ and its complementary space, respectively. Let $\tilde{Q}_{T_{0}}=I_{n_{0}}$ and $\tilde{Q}$ be the diagonal block matrix of $\{\tilde{Q}_{T_{k}}\}_{k=0}^{K}$. We define the transformed design matrix as $\tilde{X}^{(k)}(T_{k})=\tilde{Q}_{T_{k}}^{1/2}X^{(k)}$ and the stacked transformed design matrix as $\tilde{\boldsymbol{X}}(T)=({X}^{(0)\top},\tilde{X}^{(1)}(T_{1})^{\top},\ldots,\tilde{X}^{(K)}(T_{K})^{\top})^{\top}\in\mathbb{R}^{N\times p}$. We further define $\Phi_{S,T}=\tilde{\boldsymbol{X}}_{S}(T)(\tilde{\boldsymbol{X}}_{S}(T)^{\top}\tilde{\boldsymbol{X}}_{S}(T))^{-1}\tilde{\boldsymbol{X}}_{S}(T)^{\top}$.
Given a model $\{S,T\}$, we define the covariate-$j$-specific transferable set as $H^{(j)}_{S,T}=\{k\in[K]:j\in S\setminus T_{k}\}\cup\{0\}$. Denote $H_{S,T}=\cup_{j\in S}H_{S,T}^{(j)}$ as the index set of all least-transferable sources given $\{S,T\}$, indicating that all sources in $H_{S,T}$ have at least one covariate similar to the target. We further define $n^{(j)}_{S,T}=\sum_{k\in H^{(j)}_{S,T}}n_{k}$ to be the useful sample size for covariate $j$ and $n_{S,T}=\sum_{k\in H_{S,T}}n_{k}$ as the sample size of all least-transferable sources given model $\{S,T\}$. Similarly to $\tilde{\boldsymbol{X}}(T)$ defined above, we use $\tilde{\boldsymbol{X}}_{H_{S,T}}$ to denote the stacked design matrix consisting of all the least-transferable sources in $H_{S,T}$.
4.1 Oracle theoretical results with known $T^{*}$
We first consider the oracle setting where the similarity structures $\{T^{*}_{k}\}_{k=1}^{K}$ of the sources are known in advance. For ease of notation, when there is no ambiguity, we omit the dependence of $\{{\boldsymbol{X}}(T^{*}),\Phi_{S,T^{*}},H_{S,T^{*}},n^{(j)}_{S,T^{*}},n_{S,T^{*}}\}$ on $T^{*}$ in this subsection, and denote them as $\{\boldsymbol{X},\Phi_{S},H_{S},n^{(j)}_{S},n_{S}\}$, respectively. Here are the main assumptions.
Assumption 1 (True model).
[…]
Assumptions 1(i)-(ii) require that the order of the total signal strength of each dataset is no more than $n_{k}\log p$ and the non-influential signal strength in the transformed dataset is not large (Yang et al., 2016). Assumption 1(iii) requires that the transferable sample size along each dimension is comparable to the total sample size, excluding true models where only a small portion of sources are useful along each dimension.
Remark 2.
The definition of $T_{k}^{*}$ in (5) requires that $\|\boldsymbol{\delta}^{(k)}_{T_{k}^{*\mathrm{c}}}\|_{2}=0$. With more precise analysis, we allow for approximately similar sources with $\|\boldsymbol{\delta}^{(k)}_{T_{k}^{*\mathrm{c}}}\|^{2}_{2}\lesssim(\sigma^{(0)}_{y})^{2}\log p/n_{k}$, which aligns with the spirit of Assumption 1(ii). We assume that they are zero for simplicity.
Assumption 2 (Design matrix).
(i) (Normalized columns) The design matrix on source $k$ has been normalized so that $\|{{X}}^{(k)}_{j}\|_{2}^{2}=n_{k}$ for all $j\in[p]$ and $k\in[K]\cup\{0\}$;
(ii) (Restricted eigenvalue) The gram matrix of the considered model space satisfies that $\min_{\{S:|S|\leq L_{s}\}}\lambda_{\min}({\tilde{\boldsymbol{X}}_{H_{S},S}^{\top}\tilde{\boldsymbol{X}}_{H_{S},S}}/{n_{S}})\geq\lambda_{l}$ for some $\lambda_{l}\in(0,1)$, $\lambda_{\operatorname{min}}(X^{(0)\top}_{S^{*}}X^{(0)}_{S^{*}})\geq n_{0}\lambda_{l}$, and $\lambda_{\operatorname{min}}(X^{(k)\top}_{T_{k}^{*}}X^{(k)}_{T_{k}^{*}})\geq n_{k}\lambda_{l}$ for all $k\in[K]$ with $|T_{k}^{*}|\geq 1$;
(iii) (Sparse projection) Let ${Z}\sim\mathcal{N}(0,{I}_{N})$. For some $C_{3}\geq 8{\lambda_{l}}^{-1}$, it holds that
$\mathbb{E}_{Z}\left[\max_{S_{1},S_{2}:|S_{1}|\leq|S_{2}|\leq L_{s}}\max_{j\in S_{2}\setminus S_{1}}\frac{1}{\sqrt{n_{S_{2}}^{(j)}}}\left|{Z}^{T}\tilde{Q}(I_{N}-\Phi_{S_{1}})\tilde{\boldsymbol{X}}_{j}\right|\right]\leq\frac{\sqrt{C_{3}\lambda_{l}\log p}}{2}.$
(iv) (Weak correlation) For all $k\in[K]$, and $j\in[p]$, the transformed covariates satisfy that $\|\tilde{X}_{j}^{(k)}\|^{2}_{2}\geq d_{1}n_{k}$ and $\tilde{\boldsymbol{X}}_{j}^{\top}(I-\Phi_{S})\tilde{\boldsymbol{X}}_{j}\geq d_{1}d_{2}n^{(j)}_{S}$ for all $S$ satisfying $|S|\leq L_{s}$, with some $d_{1},d_{2}\in(0,1)$.
Assumption 2 imposes conditions on the projected design, which are parallel to the assumptions for single-task high-dimensional regression in the Bayesian literature (Narisetty and He, 2014; Yang et al., 2016). Assumption 2(ii) allows the model size $S$ to grow slowly compared to $p$ and is satisfied when $L_{s}\leq\min\{p,n/\log p\}$ and Assumption 2(iv) holds. Assumptions 2(iii) and 2(iv) are mild and can hold with $C_{3}=4\lambda_{l}^{-1}$ and $d_{1}=d_{2}=1$ under orthogonal design.
Assumption 3 (Prior).
Let $\underline{\tau}=\min_{k}|{\tau}_{k}|$. Assume that the prior satisfies (i) $q_{0}\leq p^{-c}$ with $c>2+{3}C_{1}{\lambda_{l}}^{-1}(\eta^{-2}+{\underline{\tau}}^{-2})/2+C_{2}+C_{3}$; (ii) $\max_{j\in S^{*}}n^{(j)}_{S^{*}}\eta^{2}\leq Cp^{2r}$ for some $r>0$; (iii) $n_{k}\tau_{k}^{2}\leq Cp^{2r_{k}}$ for some $r_{k}>0$ for $k\in[K]$.
Assumption 3 is on the choice of prior hyperparameters, which is similarly assumed in Narisetty and He (2014) and Narisetty et al. (2019).
Assumption 4 (Large signal).
The signal strength in the target significant set $S^{*}$ satisfies
$\zeta_{j}^{2}\geq C_{\beta}\frac{(\sigma_{y}^{(0)})^{2}\log p}{n^{(j)}_{S^{*},T^{*}}d_{1}d_{2}},\text{ for all }j\in S^{*},$
with constant $C_{\beta}=8({3}C_{1}{\lambda_{l}}^{-1}(\eta^{-2}+{\underline{\tau}}^{-2})/2+r+c+2)+4(C_{2}+C_{3})$.
Assumption 4 is parallel to the $\beta$-min condition in the high-dimensional literature for variable selection consistency (Bühlmann and Van De Geer, 2011), which requires the signal strength on the significant set to have a large enough magnitude. Note that the minimal signal strength restriction here is covariate-specific, depending on how many transferable sources there are along with the specific dimension.
Theorem 1 (Model selection).
Suppose $25\max\{1,s^{*}\}\log p\leq\min_{\{S:|S|\leq L_{s}\}}n_{S}$. Under Assumptions 1- 4, the posterior of the true model $S^{*}$ satisfies that,
$\Pi(S^{*}|\{y^{(k)}\}_{k=0}^{K})\geq 1-\tilde{c}_{1}p^{-1},$
for some constant $\tilde{c}_{1}>0$.
Theorem 1 can be regarded as an extension of existing results for single-task Bayesian high-dimensional regression (Narisetty and He, 2014; Yang et al., 2016) to the domain of transfer learning. We relax the conditions for single-task Bayesian high-dimensional regression from two perspectives.
First, compared to the requirements outlined in Yang et al. (2016) and Narisetty et al. (2019), where a minimum signal strength constraint of $\min_{j\in S^{*}}(\beta^{(0)*}_{j})^{2}\gtrsim\log p/n_{0}$ is necessary, we relax it in Assumption 4 by replacing the target-only sample size $n_{0}$ in the denominator with a covariate-$j$-specific transferable size $n_{S,T^{*}}^{(j)}$ of all the sources and target. As the number of similar sources along the $j$-th covariate becomes larger, we can detect $\beta_{j}^{(0)*}$ more effectively. In the ideal scenario of $j\in\cap_{k=1}^{K}T_{k}^{*\mathrm{c}}$, we have $n_{S^{*},T^{*}}^{(j)}=N$, where sources are fully utilized. In the worst case where all sources are non-transferable on $S$, i.e., $j\in\cap_{k=1}^{K}T_{k}$, $n_{S^{*},T^{*}}^{(j)}$ remains at $n_{0}$, resembling the case where only target data is utilized.
Second, we allow the dimension $p$ of the covariates to be higher compared to the restriction $\max\{1,s^{*}\}\log p\lesssim n_{0}$ required in (Yang et al., 2016). When all sources are mostly informative, we obtain $\min_{\{S:|S|\leq L_{s}\}}n_{S}=N$, which allows $\log p$ to go as large as the total sample size. Even in the worst case with no informative sources, $\min_{\{S:|S|\leq L_{s}\}}n_{S}$ remains at $n_{0}$, showing the robustness of our method.
Based on the theoretical results on model selection in Theorem 1, we can establish the posterior contraction rates of our proposed method in $\ell_{1}$ loss and $\ell_{2}$ loss to guarantee parameter estimation.
Theorem 2 (Estimation).
Under the assumptions in Theorem 1 and Assumption LABEL:assump:eigen in the supplementary material, the posterior satisfies that,
$\Pi\left(\boldsymbol{\beta}^{(0)}:\|\boldsymbol{\beta}^{(0)}-\boldsymbol{\beta}^{(0)*}\|_{1}\geq M\sigma^{(0)}_{y}s^{*}\sqrt{\frac{\log p}{n_{S^{*}}}}~{}\bigg{|}~{}\{y^{(k)}\}_{k=0}^{K}\right)\leq\tilde{c}_{2}p^{-1},$
$\Pi\left(\boldsymbol{\beta}^{(0)}:\|\boldsymbol{\beta}^{(0)}-\boldsymbol{\beta}^{(0)*}\|_{2}\geq M\sigma^{(0)}_{y}\sqrt{\frac{s^{*}\log p}{n_{S^{*}}}}~{}\bigg{|}~{}\{y^{(k)}\}_{k=0}^{K}\right)\leq\tilde{c}_{2}p^{-1},$
for some constant $\tilde{c}_{2},M>0$.
Theorem 2 indicates that the posterior allocates most of its mass in a neighborhood around the true parameter in terms of $\ell_{1}$ and $\ell_{2}$ loss. The size of the neighborhood depends on $n_{S^{*}}$ indicating that the contraction rates are faster than those when only target data is utilized, thus leading to better estimation behaviors. The $\ell_{1}$ and $\ell_{2}$ error convergences of our method hold under weaker conditions of $(s^{*})^{2}\log p=o(n_{S^{*}})$ and $s^{*}\log p=o(n_{S^{*}})$ than the conventional single-task conditions of $(s^{*})^{2}\log p=o(n_{0})$ and $s^{*}\log p=o(n_{0})$, respectively. Compared with the existing theoretical results of transfer learning (Li et al., 2022; Tian and Feng, 2023; Li et al., 2023), we are able to obtain better error bounds when global similarity is not satisfied.
We can further prove the posterior contraction rate of our proposed method in terms of prediction error under regularity assumptions.
Theorem 3 (Prediction).
Under the assumptions in Theorem 1 and Assumption LABEL:assump:eigen-LABEL:assump:pred in the supplementary material, the posterior satisfies that,
$\Pi\left(\boldsymbol{\beta}^{(0)}:\|X^{(0)}(\boldsymbol{\beta}^{(0)}-\boldsymbol{\beta}^{(0)*})\|_{2}\geq M_{1}\sigma^{(0)}_{y}s^{*}\sqrt{\frac{n_{0}\log p}{n_{S^{*}}}}~{}\bigg{|}~{}\{y^{(k)}\}_{k=0}^{K}\right)\leq\tilde{c}_{3}p^{-1},$
for some constant $\tilde{c}_{3},M_{1}>0$.
Similar to Theorem 2, Theorem 3 indicates most of the posterior probability concentrates in the ball around the truth in terms of the prediction error with a faster contraction rate than the single-task Bayesian regression (Castillo et al., 2015).
The main difference between Theorems 1-3 and existing theories of transfer learning (Li et al., 2022; Tian and Feng, 2023; Li et al., 2023) lies in the fact that the benefits of transfer learning brought by our estimator are through each variable rather than the entire set of parameters. Therefore, when there is significant heterogeneity in the source datasets, especially when some variables differ significantly from those in the target dataset, existing methods may fail to extract useful information from the sources due to large overall parameter differences. In contrast, our method can still extract information from similar parts of heterogeneous sources and integrate partial information from each source data to jointly improve the estimation of the target parameters.
4.2 Adaptive theoretical results with unknown $T^{*}$
In practice, the non-transferable sets $\{T^{*}_{k}\}_{k=1}^{K}$ are usually unknown. Therefore, more restrictive assumptions are needed for the simultaneous detection of the significant target signal set and the source-specific non-transferable sets. We provide the following strengthened assumptions.
Assumption 2S (Design matrix).
(i) (Orthogonal design) For $k\in[K]\cup\{0\}$, $\|{{X}}^{(k)}_{j}\|_{2}^{2}=n_{k}$ for all $j\in[p]$ and ${{X}}^{(k)\top}_{j_{1}}{{X}}^{(k)}_{j_{2}}=0$ for all $j_{1}\neq j_{2},j_{1},j_{2}\in[p]$;
(ii) (Restricted eigenvalue) The gram matrix of the considered model space satisfies that $\min_{\{S,T:|S|\leq L_{s},\max_{k}|T_{k}|\leq L_{t}\}}\lambda_{\min}({\tilde{\boldsymbol{X}}_{H_{S},S}(T)^{\top}\tilde{\boldsymbol{X}}_{H_{S},S}(T)}/{n_{S,T}})\geq\lambda_{l}$ for some $\lambda_{l}\in(0,1)$;
(iii) (Sparse projection) Let ${Z}\sim\mathcal{N}(0,{I}_{N})$. For some $C_{3}\geq 8{\lambda_{l}}^{-1}$, it holds that
$\mathbb{E}_{Z}\left[\max_{S_{1},S_{2}:|S_{1}|\leq|S_{2}|\leq L_{s}}\max_{j\in S_{2}\setminus S_{1}}\max_{T:\max_{k}|T_{k}|\leq L_{t}}\frac{1}{\sqrt{n_{S_{2}}^{(j)}}}\left|{Z}^{T}\tilde{Q}(I_{N}-\Phi_{S_{1},T})\tilde{\boldsymbol{X}}_{j}(T)\right|\right]\leq\frac{\sqrt{C_{3}\lambda_{l}\log p}}{2};$
(iv) (Sparse projection on source $k$) Let ${Z}_{k}\sim\mathcal{N}(0,{I}_{n_{k}})$, $k\in[K]$. For some $C_{3}^{(k)}\geq 8{\lambda_{l}}^{-1}$,
$\mathbb{E}_{Z}\left[\max_{T_{k},\tilde{T}_{k}:|T_{k}|\leq|\tilde{T}_{k}|\leq L_{t}}\max_{j\in\tilde{T}_{k}\setminus{T}_{k}}\left|{Z}_{k}^{T}\tilde{X}^{(k)}_{j}(T_{k})\right|\right]\leq{\sqrt{C_{3}^{(k)}\lambda_{l}\log p}}/{2}.$
Assumption 2S modifies Assumption 2 to account for the case with unknown $T^{*}$. We adopt the orthogonal design Assumption 2S(i) here for simplicity in notation and theoretical analysis. This assumption could be relaxed to a weak correlation assumption, similar to Assumption 2(iv), but with a more refined analysis. Assumption 2S(ii) implicitly considers the model space $\{S,T:\min_{j\in[p]}n_{S,T}^{(j)}\geq c^{\prime}_{0}n_{S,T},|S|\leq L_{s},\max_{k\in[K]}|T_{k}|\leq L_{t}\}$ for some constant $c^{\prime}_{0}>0$. Assumptions 2S(iii) and 2S(iv) are mild and hold automatically with $C_{3}=C_{3}^{(k)}=4\lambda_{l}^{-1}$ under the orthogonal design.
Assumption 3S (Prior).
Assume that the prior satisfies (i) Assumption 3; (ii) $q_{k}\leq p^{-c_{k}}$ with $q_{k}\leq p^{-c_{k}}$ with $c_{k}>2+C_{3}^{(k)}+{3}C_{1}{\lambda_{l}}^{-1}{\underline{\tau}}^{-2}/{2}$ for $k\in[K]$.
In addition to Assumption 3, Assumption 3S(ii) imposes additional assumptions on the prior for the similarity of each source to the target.
Assumption 4S (Large signal).
For all candidate models $T$ satisfying Assumption 2S, assume that the signal strength of covariate $j$ in the target significant set $S^{*}$ satisfies
$\zeta_{j}^{2}\geq C_{\beta}\frac{(\sigma_{y}^{(0)})^{2}\log p}{{n}^{(j)}_{S^{*},T}},$
with constant $C_{\beta}=8({3}C_{1}{\lambda_{l}}^{-1}(\eta^{-2}+{\underline{\tau}}^{-2})/{2}+r+c+2)+4(C_{2}+C_{3})$.
Assumption 4S replaces the ideal sample size $n^{(j)}_{S^{*},T^{*}}$ in the denominator of Assumption 4 with an adaptive ${n}^{(j)}_{S^{*},T}$, which can be seen as a cost of searching for the unknown non-transferable set $T^{*}$. It is important to note that Assumption 4S is no more restrictive than the minimal signal strength of at least the order $\sqrt{\log p/n_{0}}$ needed when using the target data only. This is because $n_{S^{*},T}^{(j)}\geq n_{0}$ always holds, regardless of the model $T$ we consider.
Assumption 5.
Denote $\tilde{n}^{(j)}_{S^{*},T^{*}\setminus T}=\sum_{\{k\in[K]:j\in S^{*}\cap(T_{k}^{*}\setminus T_{k})\}}n_{k}$. Assume that
$\min_{T:\sum_{k}|T_{k}^{*}\setminus T_{k}|\geq 1,\max_{k}|T_{k}|\leq L_{t}}\min_{j\in S^{*}\cap(\cup_{k=1}^{K}T_{k}^{*}\setminus T_{k})}\left(1-\frac{\tilde{n}^{(j)}_{S^{*},T^{*}\setminus T}}{n^{(j)}_{S^{*},T}}\right)\geq C_{l}.$
Assumption 6 (Large dissimilarity).
We assume that the minimal dissimilarity on non-transferable set $T_{k}^{*}$ of source $k$ satisfies
$\min_{j\in T_{k}^{*}}(\delta_{j}^{(k)*})^{2}\geq C_{\delta}\frac{(\sigma_{y}^{(0)})^{2}\log p}{n_{k}},$
for some $C_{\delta}>2C_{l}^{-1}(c_{k}+3C_{1}{{\lambda}_{l}}^{-1}{\underline{\tau}}^{-2}/2+C_{2}/2)+2$.
Assumption 5 and 6 ensure that the probability of missing elements of true model $T^{*}$ tends to zero. Note that in Assumption 5, $n^{(j)}_{S^{*},T}$ is always strictly greater than $\tilde{n}_{S,T^{*}\setminus T}^{(j)}$ as the target set is always included in $H_{S^{*},T}^{(j)}$. Assumption 6 requires the minimal difference strength on the non-transferable signals of source $k$ to be at least the order of $\sqrt{n_{k}/\log p}$ to ensure source dissimilarity detection consistency, which is in the same spirit of the minimal signal strength restriction in Assumption 4S for target signal detection consistency.
Theorem 4 (Model selection).
Suppose $25\max\{1,s^{*}\}\log p\leq\min_{\{S:|S|\leq L_{s}\}}n_{S,T^{*}}$. Under Assumptions 1, 2S-4S, 5, and 6, the posterior of $S^{*},T^{*}$ satisfies that,
$\Pi(S^{*},T^{*}|\{y^{(k)}\}_{k=0}^{K})\geq 1-\tilde{c}_{1}p^{-1},$
for some constant $\tilde{c}_{1}>0$. Consequently, we have the covariate-specific source selection consistency, that is,
$\Pi(S^{*},\{H_{S^{*},T^{*}}^{(j)}\}_{j=1}^{p}|\{y^{(k)}\}_{k=0}^{K})\leq 1-\tilde{c}_{1}p^{-1}.$
Theorem 4 establishes both the target signal detection consistency and the covariate-specific source detection consistency simultaneously, representing a robust and adaptive version of the oracle results in Theorem 1. Regarding robustness, the CONCERT method performs at least as well as the target-only method. For adaptivity, note that the model space of $T$ we search is much smaller in practice. When we restrict our interest to the model space $\mathcal{M}_{T}:=\{T:\min_{j\in[p]}n_{[p],T}^{(j)}\geq c^{\prime}_{0}n_{0}\}$ for some $c^{\prime}_{0}>1$, we can still relax the minimal signal strength requirement in Assumption 4S and detect weaker signals.
Theorem 5 (Estimation).
Under the assumptions in Theorem 4 and Assumption LABEL:assump:eigen in the supplementary material, the posterior satisfies that,
$\Pi\left(\boldsymbol{\beta}^{(0)}:\|\boldsymbol{\beta}^{(0)}-\boldsymbol{\beta}^{(0)*}\|_{1}\geq M\sigma^{(0)}_{y}s^{*}\sqrt{\frac{\log p}{n_{S^{*}}}}~{}\bigg{|}~{}\{y^{(k)}\}_{k=0}^{K}\right)\leq\tilde{c}_{2}p^{-1},$
$\Pi\left(\boldsymbol{\beta}^{(0)}:\|\boldsymbol{\beta}^{(0)}-\boldsymbol{\beta}^{(0)*}\|_{2}\geq M\sigma^{(0)}_{y}\sqrt{\frac{s^{*}\log p}{n_{S^{*}}}}~{}\bigg{|}~{}\{y^{(k)}\}_{k=0}^{K}\right)\leq\tilde{c}_{2}p^{-1},$
for some constant $\tilde{c}_{2},M>0$.
Theorem 6 (Prediction).
Under the assumptions in Theorem 4 and Assumption LABEL:assump:eigen-LABEL:assump:pred in the supplementary material, the posterior satisfies that,
$\Pi\left(\boldsymbol{\beta}^{(0)}:\|X^{(0)}(\boldsymbol{\beta}^{(0)}-\boldsymbol{\beta}^{(0)*})\|_{2}\geq M_{1}\sigma^{(0)}_{y}s^{*}\sqrt{\frac{n_{0}\log p}{n_{S^{*}}}}~{}\bigg{|}~{}\{y^{(k)}\}_{k=0}^{K}\right)\leq\tilde{c}_{3}p^{-1},$
for some constant $\tilde{c}_{3},M_{1}>0$.
Theorem 5 and 6 show that the CONCERT guarantees preferable posterior contraction rates in terms of $\ell_{1}/\ell_{2}$ loss and prediction error under the unknown $T^{*}$ case, which extends the results in Theorem 2 and 3.
5 Simulation Studies
In this section, we empirically evaluate the performance of our proposed method CONCERT and compare it with various baseline target methods and state-of-art transfer learning methods in extensive simulation studies. Specifically, target methods include Lasso (Tibshirani, 1996), SparseVB with Laplace slab prior (Ray and Szabó, 2022), and NaiveVB where CONCERT simplifies to a solo of target data. Transfer methods include TransLasso (Li et al., 2022) and TransGLM (Tian and Feng, 2023).
We consider both linear and logistic regression and present the estimation errors in three different settings. In Section 5.1, we examine the performance of all methods in the framework of informative set, a scenario commonly adopted in the transfer learning literature. In Section 5.2, we consider the case where sources are heterogeneous and only share partial information with the target. In Section 5.3, all sources share the whole information with the target but with redundant signals.
We also compare the prediction errors and variable selection results in Section LABEL:append:prediction and LABEL:append:selection in Supplementary Material, respectively. The posterior similarity probabilities of CONCERT are investigated in Section LABEL:append:source_var. We study the sensitivity of our CONCERT to hyperparameters and compare the run time with various methods (including the MCMC alternative) in Section LABEL:append:hyper and LABEL:append:time, respectively. Simulations under other settings are conducted in Section LABEL:append:simu.
5.1 Existence of Informative Set
Motivated by Li et al. (2022), we first consider the scenario where there exists a so-called informative set $\cal A$. If a source $k\in\cal A$, the similarity between this source and the target is high in the sense that the contrast $\boldsymbol{\delta}^{(k)}$ enjoys a small norm and sparse structure. In contrast, if a source $k\not\in{\cal A}$, i.e., is non-informative, its parameter deviates from the target parameter more severely. Note that such non-informativeness is based on a global consideration, while locally, potential covariates in the source data may be helpful.
We take $n_{0}=150$, $n_{k}=100$, for $k=1,\ldots,10$ and $p=200$. The covariates $x^{(k)}_{i}$ are generated independently from $\mathcal{N}(0,1)$. We set $\boldsymbol{\beta}^{(0)}=(\mathbf{0.5}^{\top}_{s},\mathbf{0}^{\top}_{p-s})^{\top}$ for linear regression and $\boldsymbol{\beta}^{(0)}=(\mathbf{1}^{\top}_{s},\mathbf{0}^{\top}_{p-s})^{\top}$ for logistic regression with $s=16$. Given $\cal A$, if $k\in{\cal A}$, we set
$\beta^{(k)}_{j}=\left\{\begin{array}[]{ll}\beta_{j}^{(0)}+0.5\operatorname{Sign}(j),&j\in H_{k};\\ \beta_{j}^{(0)},&\text{otherwise,}\end{array}\right.$
where $H_{k}$ is a random index subset of $[p]$ with $|H_{k}|=h$ and $\operatorname{Sign}(j)$ takes ±$\pm 1$ with equal probability. If $k\not\in{\cal A}$, we set $\beta^{(k)}_{j}=\beta_{j}^{(0)}+\operatorname{Sign}(j)$ if $j\in H_{k}$ with $|H_{k}|=20$. We consider varying $|{\cal A}|$ and $h$ with $|{\cal A}|\in\{0,2,4,6,8,10\}$ and $h\in\{4,8,12\}$. We run the experiment on 100 independent trials and summarize the means and standard errors of the rooted sum of squared estimation errors of $\boldsymbol{\beta}^{(0)}$ for each method. The results are reported in Figure 2.
Figure 2: Estimation errors with different informative set numbers $|A|$ and varying sizes $h$.
As we observe from Figure 2, CONCERT exhibits a significant advantage over all other candidates. We underscore three highlights. First, our CONCERT significantly improves the performance over traditional VB estimators, indicating the benefits of transfer learning. Though target VB methods perform worse than Lasso in the logistic setting, CONCERT outperforms TransGLM considerably. Second, all the transfer learning methods perform worse as the size of informative sources decreases. However, CONCERT is much less sensitive to the decrease of globally informative sources, especially in the linear setting. When all the sources are globally non-informative, CONCERT maintains excellent performance, while other transfer learning methods deteriorate to target Lasso or even show negative transfer. This is because our covariate-specific informative indicators $I_{j}^{(k)}$ can adaptively include locally informative covariates in different sources. Third, CONCERT is also robust to the varying size $h$, which is also due to our partial similarity modeling.
5.2 Heterogeneous Information
In this subsection we consider a more general scenario where sources contain heterogeneous information and only share some common signals with the target. In the sense of informative set, they should be considered non-informative.
Here the covariates and noises are generated in the same manner as that in Section 5.1. For the parameters, we set $\boldsymbol{\beta}^{(0)}=(\mathbf{0.5}^{\top}_{s},\mathbf{0}^{\top}_{p-s})^{\top}$ for linear regression and $\boldsymbol{\beta}^{(0)}=(\mathbf{1}^{\top}_{s},\mathbf{0}^{\top}_{p-s})^{\top}$ for logistic regression with varying $s\in\{4,8,12,16\}$, and each $\boldsymbol{\beta}^{(k)}$ only shares partial values of $\boldsymbol{\beta}^{(0)}$. Specifically, we define the informative signal ratio $w$ as the ratio of shared target signals between $\boldsymbol{\beta}^{(k)}$ and $\boldsymbol{\beta}^{(0)}$ among the first $s$ components. Then we set
$\beta^{(k)}_{j}=\left\{\begin{array}[]{ll}0,&j\in[s]/W_{k};\\ \beta_{j}^{(0)}+0.5\operatorname{Sign}(j),&j\in U_{k};\\ \beta_{j}^{(0)},&\text{otherwise,}\end{array}\right.$
where $W_{k}$ is a random subset of $[s]$ with $|W_{k}|=\lceil ws\rceil$ and $U_{k}$ is a random subset of $\{s+1,\ldots,p\}$ with $|U_{k}|=20$. We set the ratio to take values in $\{0.6,0.7,0.8,0.9,1\}$ and present the results based on 100 replications in Figures 3.
Figure 3: Estimation errors with different informative signal ratios and target signal sizes.
From Figure 3, it can be seen that the challenge of information transfer intensifies as the target information becomes more complicated and the source heterogeneity increases. However, CONCERT can still improve much compared with NaiveVB and SparseVB. We highlight two points here. Firstly, CONCERT outperforms other transfer methods significantly. As in most cases, all sources are globally non-informative, TransLasso and TransGLM fail to extract useful knowledge from source data, which makes them fail to compete with Lasso on target data. However, our CONCERT succeeds in identifying locally informative covariates in most cases and shows strong robustness when the amount of shared information is not dominant. Secondly, though the performance of all transfer methods improves as the informative signal ratio becomes larger, CONCERT is much more responsive. We also note that the advantage of CONCERT over other candidates becomes much clearer as the sparsity number $s$ increases.
5.3 Redundant Information
We further investigate the power of CONCERT on information transfer under the scenario where all sources contain complete target signals, however, with various degrees of contamination on the rest. Though intrinsically these sources are all excellent candidates for transfer learning, many of them are regarded as globally non-informative and can be removed under the framework of informative set.
We adopt the same data-generating process for covariates and noises as before. We set the target parameter as $\boldsymbol{\beta}^{(0)}=(\mathbf{0.5}^{\top}_{s},\mathbf{0}^{\top}_{p-s})^{\top}$ for linear regression and $\boldsymbol{\beta}^{(0)}=(\mathbf{1}^{\top}_{s},\mathbf{0}^{\top}_{p-s})^{\top}$ for logistic regression with $s=16$ and the source parameters are constructed according to
$\beta^{(k)}_{j}=\left\{\begin{array}[]{ll}\beta_{j}^{(0)}+\rho\operatorname{Sign}(j),&j\in U_{k};\\ \beta_{j}^{(0)},&\text{otherwise,}\end{array}\right.$
where $\rho$ represents the redundant signal strength with $\rho\in\{0.3,0.5,1,1.5\}$ and $|U_{k}|\in\{4,8,12,16,20\}$ is the redundant signal number. We repeat the experiment 100 times and the summarized results are shown in Figure 4.
Figure 4: Estimation errors with different redundant signal numbers and strengths $\rho$.
Figure 4 shows that CONCERT can avoid being misled by redundant signals and performs the best. While other transfer methods deteriorate rapidly as the redundant signal numbers and strengths grow, CONCERT can identify locally informative covariates, which reduces its bias even when the magnitude of redundant signals dominates that of true signals. We want to highlight that other transfer methods may suffer from negative transfer, especially in the cases of large redundant signal strengths, while CONCERT remains robust in all cases and is generally not affected by redundant signals in linear regression.
6 Real Data Analysis
In this section, we apply our CONCERT to GTEx data and Lending Club data to demonstrate its effectiveness under the linear and logistic regression settings, respectively.
6.1 Application to GTEx Data
The proposed transfer learning algorithm is applied to the Genotype-Tissue Expression (GTEx) data, accessible at https://gtexportal.org/. This dataset encompasses gene expression levels from 49 tissues, involving 838 individuals and aggregating 1,207,976 observations of 38,187 genes. Building upon the methodology outlined by Li et al. (2022), our investigation focuses on gene regulations within the central nervous system (CNS) across diverse tissues. The collection of CNS-related genes is consolidated under MODULE_137, comprising a total of 545 genes, alongside an additional 1,632 genes that exhibit significant enrichment in the same experiments as those within the module.
Our specific interest lies in the prediction of the expression level of the gene SH2D2A (SH2 Domain Containing 2A) in a target tissue through the utilization of other central nervous system genes. SH2D2A encodes a T-cell-specific adapter protein that regulates early T-cell activation. It plays a crucial role in inflammatory neuropathies, particularly chronic inflammatory demyelinating polyneuropathy (Uncini et al., 2011). Six brain tissues are considered target tissues and, for each target, the remaining 48 tissues serve as sources. The corresponding models are individually estimated. We first screen the genes by examining their correlation with the expression level of SH2D2A, leaving 149 genes whose correlation coefficients are greater than 0.25. All the expression levels are standardized. The average sample size for the target dataset is 194, while for the sources, it is 17,135. We randomly split the data into a training set (80% of all samples) and a testing set (20%) 50 times, and compare the performance of CONCERT with the same candidates as listed in Section 5. The average prediction error on the testing sets is reported in Table 1.
Table 1: Prediction errors (standard errors) over 50 replications on 6 different brain tissues.
Method A.C.Cortex Amygdala Cortex Hippocampus N.A.B.Ganglia S.Nigra
CONCERT 0.214 (0.046) 0.151 (0.065) 0.617 (0.092) 0.356 (0.064) 0.529 (0.166) 0.690 (0.124)
SparseVB 0.449 (0.354) 0.232 (0.202) 0.830 (0.249) 0.426 (0.143) 0.709 (0.238) 0.821 (0.161)
NaiveVB 0.279 (0.108) 0.163 (0.085) 0.820 (0.236) 0.441 (0.155) 0.633 (0.178) 0.806 (0.157)
Lasso 0.544 (1.033) 0.176 (0.134) 0.725 (0.186) 0.460 (0.286) 0.613 (0.198) 0.754 (0.176)
TransLasso 0.235 (0.078) 0.172 (0.049) 0.679 (0.160) 0.424 (0.210) 0.564 (0.207) 0.717 (0.133)
TransGLM 0.245 (0.153) 0.164 (0.075) 0.705 (0.178) 0.430 (0.127) 0.565 (0.172) 0.762 (0.141)
In general, our CONCERT performs the best among all the methods. We underscore three findings. Firstly, in all studies CONCERT makes improvements over SparseVB and NaiveVB which solely use target information, indicating that at least one of the sources is partially informative. Secondly, CONCERT outperforms all other transfer methods and enjoys a lower standard deviation. Specifically, for the Hippocampus tissue, CONCERT reduces the prediction error of Lasso by more than 20%, whereas TransLasso and TransGLM only reduce it by less than 10%. This suggests that the source datasets for Hippocampus may have limited useful global information, but there are many local pieces of information that are valuable. Such partial information can be captured by CONCERT, thereby further reducing prediction errors.
Lastly, we observe negative transfer in the Nigra tissue when using TransGLM. This may be due to that the source detection method in TransGLM is based on the global similarity gap assumption, which may not hold here. In contrast, the CONCERT remains robust and provides an improvement over target-only methods.
6.2 Application to Lending Club Data
In this section, we study the issuing of loan defaults using the Lending Club data available through Kaggle. Lending Club is a financial technology company which offers peer-to-peer (P2P) lending services through its online platform, and the Lending Club dataset contains features on both borrowers and loans issued. We consider a binary response based on the loan status, where loans that are not fully paid are denoted as class 1, and otherwise denoted as class 0. Our goal is to predict the risk of loan default within the target domain using information from other source domains.
We focus on the period from 2007 to 2011. After excluding the variables whose values are missing for most records and adding the pairwise interaction terms between the remaining variables, we obtain 3,458 loan records and 190 predictors in six domains: vacation, medical, wedding, education, moving, and house. We consider all domains as our targets and estimate the corresponding models individually. For each target, all other domains are utilized as sources. Similarly to Section 6.1, we randomly split the data into a training set and a testing set 100 times, and compare the CONCERT with the same candidates for the logistic setting in Section 5. The cutoff for binary classification is set as $0.5$, and the average prediction error on the testing sets is reported in Table 2.
Table 2: Prediction errors (standard errors) over 50 replications on 6 different loan types.
Method Vacation Medical Wedding Education Moving House
CONCERT 0.141 (0.038) 0.173 (0.029) 0.106 (0.019) 0.208 (0.045) 0.160 (0.027) 0.167 (0.040)
SparseVB 0.154 (0.045) 0.175 (0.028) 0.108 (0.019) 0.212 (0.046) 0.165 (0.027) 0.185 (0.035)
NaiveVB 0.151 (0.040) 0.177 (0.028) 0.111 (0.021) 0.222 (0.052) 0.167 (0.029) 0.188 (0.037)
Lasso 0.388 (0.127) 0.438 (0.041) 0.379 (0.040) 0.361 (0.079) 0.595 (0.156) 0.367 (0.045)
TransGLM 0.156 (0.039) 0.174 (0.027) 0.108 (0.020) 0.213 (0.050) 0.164 (0.028) 0.178 (0.039)
Table 2 shows the advantage of the CONCERT on improving the prediction accuracy of loan default on the target domain. It is noteworthy that Bayesian methods such as SparseVB and NaiveVB can perform much better than Lasso when only the target dataset is used. Although TransGLM shows improvements over Lasso, its performance is only comparable to that of the target-only Bayesian methods. In contrast, our method can further enhance prediction accuracy of the target-only Bayesian methods. Specifically, in the domains of vacation and house, the CONCERT further reduces the prediction errors of SparseVB and NaiveVB by about 10%.
7 Conclusion
In this paper, we proposed CONCERT, a novel Bayesian method to allow partial information transfer. By imposing a conditional spike-and-slab prior together with the traditional spike-and-slab prior on the joint distribution of parameters in the sources and the target, we can achieve variable selection and information transfer simultaneously, robustly, and adaptively. Our theoretical results demonstrate the validity of CONCERT for robust partial information transfer. We make use of the variational Bayes framework to make our algorithm scalable. Numerical experiments and real data analysis show that the CONCERT is robust to source heterogeneity and can better utilize partially informative covariates from the sources to improve target learning.
There are several potential directions for future work. First, although our simulation results demonstrate that our method performs well in both linear regression and logistic regression, the current theoretical framework is limited to linear regression. Extending the theoretical properties to the generalized linear models might require additional efforts. Second, inference procedures may be developed, which is a significant advantage of Bayesian methods. Moreover, the innovative use of the conditional spike-and-slab prior is not limited to transfer learning. This idea can also be applied to other settings where both shared components and subject-specific components exist, such as data integration or robust learning. These extensions are beyond the scope of this paper and require further development.